OpenStudy (anonymous):

What are the first five terms of the pattern with the formula 21n? The first five terms are , , , , and . @TwilightRain1

OpenStudy (loser66):

21*1 =?

OpenStudy (anonymous):

1

OpenStudy (loser66):

21*2=?

OpenStudy (loser66):

21*3=?

OpenStudy (anonymous):

42

OpenStudy (loser66):

21*4=?

OpenStudy (loser66):

21*5=?

OpenStudy (anonymous):

63, 84, 105

OpenStudy (anonymous):

Thanks @lo

OpenStudy (loser66):

yeah, hence, your first five terms are {21, 42, 63,84,105,....}

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

Thanks!

OpenStudy (loser66):

easy, right? hhehehe...